What is the sum of the first 55 consecutive odd numbers
Answer by oscargut(2103)

That sum is: 1+3+5+7+......+109
= 55+ (0+2+4+6+8+....+108)
= 55+ 2(0+1+2+....+54)
= 55 + 2(54)(55)/2
= 55 + 54(55)
= (1+54)55
= 55*55
= 3025
Answer: 3025
You can ask me more at: mthman@gmail.com
Thanks